OpenStudy (anonymous):

@hoblos : Determine the slope of the line which passes through these two points: (-1, 3) and (6, 9).

OpenStudy (hoblos):

slope = (y2 - y1)/(x2 - x1) = (9-3)/(6+1) = 6/7